A company pays $30,000 for two machines. Machine A is appraised at a fair market value of $24,000 and Machine B at a fair market value of $8,000. The cost of Machine B is recorded for book purposes at 25. .. a. $7,500 b. $8,000 c. $24,000 d. $30,000 26. On July , 2016, a calendar-year corporation purchases a new passenger auto for $25,000. The maximum tax depreciation allowed on the auto in 2016 is... a. $15,000 d. $7,660 c. $11,160 For the asset in Question 26, maximum depreciation in Year 6 will be c. $1,875 Use the following information to answer Questions 28-30: On January 2, 2016, a 10-year life and a residual value of $1,000. It is depreciating the machine b. $12,500 27. a. b. $720 d. $1,617 and estimates that it will have for book purposes under the straight-line method. What is the journal entry to record Year 1 depreciation for a nonmanufacturing company? 8. a. Depreciation Expense Accumulated Depreciation-Machine 1,000 1,000 b. Inventory-Work-In-Process OH Accumulated Depreciation-Machine 1,000 1,000 c. Depreciation Expense Accumulated Depreciation-Machine 4,000 4,000 d. Inventory-work-in-Process OH Depreciation Expense 1,000 1,000
Answer #1

25. answer is option A

$7,500

The book value of machine B will taken proportionately on the basis of fair market values and total cost paid. It will be calculated as follows: ($8000*30000/(24000+8000)) = 7500

26. answer is option C

$11,160

The maximum tax depreciation on passenger automobiles in 2016 in the first year along with 50% bonus is 11160

27. answer is option C

$1875

The maximum tax depreciation on passenger automobiles in 2016 all subsequent years after the third year along with 50% bonus is 1875

28. answer is option A

Depreciation expense

1000

Accumulated depreciation - machine

1000

Depreciation = cost – residual value / estimated life = (11000-1000)/10 = $1000